OpenStudy (anonymous):

3x+2 ≤ -7 or -2x+1 ≤ 9 this is an interval problem. Would the answer be [-3,-4] or (- infinety,-3]∪ [-4,+infinety)?

OpenStudy (anonymous):

Just a sec.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Lets solve the first guy. First you can subtract 2 from both sides: \[3x \le-9\]

OpenStudy (anonymous):

so \[x \le-3\]

OpenStudy (anonymous):

xyup

OpenStudy (anonymous):

So from negative infinity, up to and including -3, this solution is satisfied.

OpenStudy (anonymous):

Since x is less than -3, it is going to the left on the number line (towards minus infinity)

OpenStudy (anonymous):

So the solution would be (-infinity,-3]

OpenStudy (anonymous):

Always remember that you use parenthesis on infinities, and since we are including -3, we can have a square bracket.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Solving the other one, we get \[x \ge -4\]

OpenStudy (anonymous):

So now we want to start at -4 on the number line, include it, and have anything greater than it be a solution (anything to the right of -4 on the number line...so out towards positive infinity)

OpenStudy (anonymous):

So its solution would be \[[-4,\infty)\]

OpenStudy (anonymous):

thanks i didn't know which way to go with it lol

OpenStudy (anonymous):

If you want to union the two, the entire solution would be (\[(-\infty,-3] \cup [-4, \infty)\]

OpenStudy (anonymous):

Assuming they are solutions to the same system

OpenStudy (anonymous):

Yup, no problem.